### v1

# WaterPark: A Robustness Assessment of Language Model Watermarking

Abstract
To mitigate the misuse of large language models (LLMs), such as disinformation, automated phishing, and academic cheating, there is a pressing need for the capability of identifying LLM-generated texts. Watermarking emerges as one promising solution: it plants statistical signals into LLMs’ generative processes and subsequently verifies whether LLMs produce given texts. Various watermarking methods (“watermarkers”) have been proposed; yet, due to the lack of unified evaluation platforms, many critical questions remain under-explored: i) What are the strengths/limitations of various watermarkers, especially their attack robustness? ii) How do various design choices impact their robustness? iii) How to optimally operate watermarkers in adversarial environments?
To fill this gap, we systematize existing LLM watermarkers and watermark removal attacks, mapping out their design spaces. We then develop WaterPark, a unified platform that integrates 10 state-of-the-art watermarkers and 12 representative attacks. More importantly, leveraging WaterPark, we conduct a comprehensive assessment of existing watermarkers, unveiling the impact of various design choices on their attack robustness. For instance, a watermarker’s resilience to increasingly intensive attacks hinges on its context dependency. We further explore the best practices to operate watermarkers in adversarial environments. For instance, using a generic detector alongside a watermark-specific detector improves the security of vulnerable watermarkers. We believe our study sheds light on current LLM watermarking techniques while WaterPark serves as a valuable testbed to facilitate future research. [^1] All the source code and data are publicly available: https://github.com/JACKPURCELL/sok-llm-watermark
1 Introduction
The recent advances in large language models (LLMs), including GPT[46], Claude[3], Mistral[22], and Llama[58], have significantly enhanced our capabilities of general-purpose text generation and complex problem-solving. However, such unprecedented capabilities also give rise to a variety of possible misuses, such as disinformation[21], automated phishing[4], and academic cheating[7]. There is thus a pressing need for the capability of identifying LLM-generated content.
A simple approach is to train classifiers that rely on the intrinsic difference between LLM- and human-generated texts[42]. Yet, with LLMs becoming increasingly capable, the gap between LLM- and human-generated texts is narrowing. Thus, watermarking is gaining attention as a promising alternative. By embedding specific statistical signals (“watermarks”) into LLMs’ generative processes, it is possible to verify whether given texts are produced by such models, as illustrated in Figure 1. A variety of LLM watermarking methods (“watermarkers") have been proposed[1, 35, 36, 26, 20, 76, 29], each with unique design choices and desirable properties, raising a set of intriguing questions:
Figure 1: LLM watermarking and watermark removal attacks.
RQ1 - What are the strengths/limitations of various watermarkers, especially their robustness against manipulations?
RQ2 - How do different design choices impact the attack robustness of watermarkers?
RQ3 - What are the best practices for operating watermarkers in adversarial environments?
Despite recent benchmarking studies on LLM watermarking, existing research is limited in addressing these questions. For instance, WaterBench[59] primarily focuses on watermarking effectiveness; MarkMyWords[49] mainly evaluates the robustness of a specific watermarker[26]; MarkLLM[47] focuses on providing a platform to compare watermark detectability, basic robustness, and text quality. Moreover, these studies lack an in-depth analysis of how a watermarker’s design choices impact its robustness. Consequently, the aforementioned questions remain largely unexplored due to the absence of unified, comprehensive evaluation platforms.
Previous Conclusion Refined Conclusion Explanation Consistency
[…]
Our Work. To bridge this critical gap, this work conducts a systematic study of state-of-the-art LLM watermarkers, focusing on their robustness against watermark removal attacks. We aim to understand how various design choices affect watermarkers’ attack resilience and identify best practices for operating watermarkers in adversarial environments. We provide insights through three key aspects:
Platform –- We develop WaterPark, the first open-source platform dedicated to evaluating the attack robustness of LLM watermarkers in a unified and comprehensive manner. As of 09/04/2024, WaterPark integrates 10 state-of-the-art watermarkers, 12 representative watermark removal attacks, and 8 key metrics. Moreover, WaterPark offers a comprehensive suite of tools for in-depth robustness assessment, including next-token distribution comparisons, attack combination analyses, and what-if scenario evaluations
Assessment –- Leveraging WaterPark, we empirically evaluate the attack resilience of representative LLM watermarkers, leading to many interesting findings, many of which challenge the conclusions in prior work, as summarized in Table 1. For instance, i) UPV is more prone to false positives than they claim, making it challenging to achieve an ideal TPR at lower FPR; ii) Some watermarks, such as UPV, may perform poorly in a single paraphrase but can still be detected after multiple paraphrases. iii) SIR claims that they can withstand paraphrase attacks, but they are unable to resist high-intensity dipper attacks or paraphrases generated by GPT. iv) UB, in addition to being resistant to token replacement attacks as mentioned in the original text, demonstrates poor robustness against other types of attacks.
Exploration – We also explore how a watermarker’s design choices impact its attack robustness, unveiling critical trade-offs between different types of robustness. For instance, i) watermarkers that rely on explicit statistical signals tend to be robust against linguistic variation attacks; ii) watermarks based on distribution transform show great resistance to attacks that mix watermarked and non-watermarked texts; iii) watermarkers’ robustness against increasingly intensive attacks largely depends on their context dependencies. Further, we also explore the best practices to operate watermarkers in adversarial environments. For instance, i) employing a generic detector alongside a watermark-specific detector improves the security of vulnerable watermarkers; ii) evaluating watermarkers’ robustness needs to consider not only individual attacks but also their combinations; iii) paraphrasing using highly capable LLMs serves as a dominant attack to test watermarkers’ robustness lower-bounds.
This study sheds light on the current LLM watermarking techniques while WaterPark servers as a useful platform to facilitate future research in this space.
2 Preliminaries
We first introduce fundamental concepts and assumptions used throughout the paper.
2.1 LLM Watermarking
[…]
At a high level, a watermark is a pattern embedded in a given signal (e.g., image and text) for identifying the signal’s source. In the context of LLMs, watermarks can be used to prove that a given text is LLM-generated (or even generated by specific LLMs). Typically, an LLM watermarking method (“watermarker”) comprises three components: the LLM, watermarking procedure (generator), and detection procedure (detector), as illustrated in Figure 1.
Typically, the generator produces watermarked texts iteratively. At each iteration, with access to a secret key $k$, the previous tokens $x_{<t}$, and the LLM’s next-token distribution $p(x_{t}|x_{<t})$, the generator generates a perturbed distribution $\tilde{p}(x_{t}|x_{<t})$, from which the next token $x_{t}$ is sampled. Meanwhile, with access to a secret key $k^{\prime}$, the detector determines whether $M$ generates a given text. Note that in symmetric schemes, $k=k^{\prime}$, while in asymmetric schemes[11], $k\neq k^{\prime}$. This study mainly focuses on symmetric watermarking schemes due to their widespread adoption.
2.2 Desiderata of Watermarking
We consider that an ideal LLM watermarker possesses the following desirable properties.
Effectiveness – The detection of (non-)watermarked texts should be highly accurate;
Fidelity – The watermarking should minimally impact the original texts’ semantics;
Robustness – The watermarked texts should be resilient against potential watermark removal attacks.
Besides, existing studies[5] also advocate indistinguishability – the watermarked and non-watermarked texts should be indistinguishable, a quality often measured by the distributional difference between the original texts and the expectation of the watermarked texts. However, recent work[49] regards this property as overly restrictive, prioritizing criteria such as effectiveness and fidelity above indistinguishability. In this study, we primarily focus on examining the robustness of current watermarking techniques and their interplay with other evaluation criteria.
2.3 Watermark Removal Attacks
One critical property of LLM watermarkers is their robustness against potential attacks. This study focuses on watermark removal attacks[25] designed to destroy the watermarks embedded in given watermarked texts while preserving their semantics. We assume a threat model similar to prior work[49, 76, 52, 27], as shown in Figure 1. The adversary intercepts the watermarked text $\tilde{T}$, modifies it, and then forwards the altered text $T^{\prime}$ to the watermark detector, aiming to avoid detection.
Adversary’s goals. The adversary modifies the watermarked text $\tilde{T}$ as $T^{\prime}$ such that the following objectives are met: effectiveness – $T^{\prime}$ evades the watermark detector (i.e., detected as non-watermarked) and quality – $T^{\prime}$ preserves $\tilde{T}$’s original semantics.
Adversary’s knowledge. Following prior work[49, 76, 52, 27], we assume the adversary has access to sample watermarked and non-watermarked texts, but cannot reproduce the watermarking procedure or interact with the detection procedure. Otherwise, it is trivial to launch adversarial attacks against the detector (e.g., by identifying and replacing green/red-listed tokens[26]).
Adversary’s capability. We assume the adversary can i) modify the watermarked text in a computationally efficient way (e.g., synonym substitution), ii) use LLMs less capable than the target LLM, and/or iii) train a detector based on given watermarked and non-watermarked texts. We argue that these assumptions are realistic and practical. Otherwise, using capable LLMs may easily generate highly-quality, non-watermarked texts to evade detection, while launching computationally expensive attacks may significantly increase the adversary’s cost.
Watermarker Information Context Dependency Generation Strategy Detection Method
Index-dep. Text-dep. Context-free Dist.-shift Dist.-reweight Dist.-transform Score-based Diff.-based Model-based
[…]
Table 2: A taxonomy of LLM watermarkers.
3 Taxonomies
We present a taxonomy of LLM watermarkers, as summarized in Table 2. The current watermarkers can be categorized based on the information carried by the watermarks (e.g., one-bit versus multi-bit) and their key design choices, including context dependency, generation strategy, and detection method. Next, we mainly focus on the key design factors.
3.1 Context Dependency
The watermarker applies a perturbation $\Delta_{t}$ to the LLM’s next-token distribution $p(x_{t}|x_{<t})$ as $\tilde{p}(x_{t}|x_{<t})=p(x_{t}|x_{<t})+\Delta_{t}$, from which the next token $x_{t}$ is generated. The perturbation $\Delta_{t}$ often depends on the given context. Three types of context dependencies are typically used in the existing watermarkers.
Index-dependent watermark. The watermarker produces a pseudo-random number $r_{t}$ by applying a keyed hash function
$r_{t}=f_{k}(t),$ (1)
that only depends on the index $t$ of the next token; $r_{t}$ is then used to generate the perturbation $\Delta_{t}$[29].
Text-dependent watermark. The watermarker considers the previous tokens $x_{<t}$ as the context window to generate $\Delta_{t}$[26, 1, 20]. For instance, GO[1] computes the hash of the concatenation of previous $w$ tokens as $r_{t}$:
$r_{t}=f_{k}(x_{t-w}\|\ldots\|x_{t-1}),$ (2)
while TGRL[26] suggests using the min hash of the previous token:
$r_{t}=\min(f_{k}(x_{t-w}),\ldots,f_{k}(x_{t-1}))$ (3)
Similarly, UB[20] concatenates the previous $w$ tokens to generate the context code for reweighting the logits. UPV[35] and SIR[36] use neural networks to generate $\Delta_{t}$ based on the previous tokens.
Context-free watermark. The watermarker applies a universal perturbation $\Delta_{t}$ across the next-token distributions of all the tokens without considering their contexts[76].
3.2 Generation Strategy
By applying the perturbation $\Delta_{t}$ to the LLM’s next-token distribution $p(x_{t}|x_{<t})$, the watermarker generates the new distribution $\tilde{p}(x_{t}|x_{<t})$ to sample the next token $x_{t}$. The existing sampling strategies can be categorized as follows.
Distribution shift. TGRL[26] modifies $p(x_{t}|x_{<t})$ by adding a shift $\delta$ to the logits of “green-list” tokens (the remaining as “red-list” tokens) as the modified distribution $\tilde{p}(x_{t}|x_{<t})$. A token $x$ is considered as green listed if $\pi_{r_{t}}(x)<\gamma d$ where $\pi_{r_{t}}$ is a permutation seeded by $r_{t}$, $\gamma$ is a parameter to control the size of the green list, and $d$ is the number of vocabulary size. Similarly, UG[76] uses a fixed red-green split over the vocabulary, showing greater robustness than TGRL against edit-distance-bounded attacks due to its “hard” split. UPV[35] selects the top-$k$ tokens from $p(x_{t}|x_{<t})$, applies a neural network to predict the green-list tokens from these tokens, and adds $\delta$ to the logits of green-list tokens. Unlike the other strategies, the distribution-shift strategy preserves the diversity of generated tokens; however, it can not be made indistinguishable, as $\tilde{p}(x_{t}|x_{<t})$ and $p(x_{t}|x_{<t})$ are inherently distinguishable.
Distribution reweight. Similar to distribution shift, this strategy alters the next-token distribution but uniquely perturbs the logit of each token. For instance, SIR[36] trains a neural network to predict the perturbation to logit of each token, which is unbiased (no preference over specific tokens) and balanced (with total perturbation summing up to 0). UB[20] advocates unbiased watermarking such that the expectation of the reweighted distribution agrees with the original distribution. It proposes two reweighting schemes: $\delta$-reweighting uniformly samples a token from $p(x_{t}|x_{<t})$ and changes its probability to 1; $\gamma$-reweighting shuffles all the tokens, rejects the first half, and double the probabilities of the remaining half. DIP[66] also uses a distribution-reweight generation strategy similar to UB but does not need to access the LM during detection.
Distribution transform. Another line of watermarkers apply randomized transform on $p(x_{t}|x_{<t})$ to sample $x_{t}$. For instance, RDF-EXP[29] and GO[1] use the Gumbel-max trick and apply the exponential transform. Let $p(x_{t}|x_{<t})=\{p_{i}\}_{i=1}^{d}$ be the distribution over the next token $x_{t}$. Then $x_{t}$ is sampled as:
$x_{t}=\arg\max_{i\leq d}r_{i}^{1/{p_{i}}}$ (4)
where $r_{i}$ is generated by the pseudo-random functions in Eq. 1 and Eq. 2. Similarly, RDF-IST[29] applies inverse transform over $p(x_{t}|x_{<t})$. With $\pi_{k}$ as a random permutation seeded by the secret key $k$, the next token $x_{t}$ is selected as:
$x_{t}=\pi_{k}\left(\min_{j\leq d}\sum_{i=1}^{j}p_{\pi_{k}(i)}\geq r_{t}\right)$ (5)
which is the smallest index in the inverse permutation such that the CDF of the next token distribution exceeds $r_{t}$. The distribution-transform strategy does not alter the next-token distribution, thus preserving the original text distribution (i.e., indistinguishability).
3.3 Detection method
The watermarker’s detector determines whether a given text $T=(x_{1},\ldots,x_{n})$ is watermarked or not. Next, we categorize the existing detection methods as follows.
Score-based detection. The detector computes the random value $r_{i}$ for each position, the per-token statistics $s(x_{i},r_{i})$, and a score over such statistics, which is then subjected to a one-tailed statistical test to determine whether the text is watermarked. The per-token statistics vary with the concrete generators. For instance, GO[1] defines $s(x_{i},r_{i})=-\log(1-h_{r_{i}}(x_{i}))$, while TGRL[26] defines $s(x_{i},r_{i})=1$ if $x_{i}$ is in the green list and 0 otherwise. One simple way to aggregate the per-token statistics is to compute their sum[1, 26, 76, 36]:
$S=\sum_{i=1}^{n}s(x_{i},r_{i})$ (6)
However, as the random values may be misaligned with the tokens (e.g., due to editing), a more robust way is to compute the alignment score (e.g., edit score[29, 60]):
$S=s^{\psi}(n,n)\;\text{s.t.}\;s^{\psi}(i,j)=\min\left\{\begin{array}[]{l}s^{\psi}(i-1,j-1)+s(x_{i},r_{j})\\ s^{\psi}(i,j-1)+\psi\\ s^{\psi}(i-1,j)+\psi\end{array}\right.$ (7)
where $\psi$ is the “edit-cost” parameter.
Differential-based detection. This line of detectors also relies on the score of a given text. However, the score is computed by comparing the given text with the non-watermarked text generated by the same LLM. For instance, UB[20] computes the log-likelihood ratio (LLR) score:
$s(i)=\log\frac{\tilde{p}(x_{i}|x_{<i})}{p(x_{i}|x_{<i})}$ (8)
and its more robust maximin variant. However, note that these detectors naturally require accessing the original LLM, which is not always feasible.
Model-assisted detection. Instead of computing the per-token statistics, which are often subject to watermark removal attacks, one may also train a model to predict whether the given text is watermarked. UPV[35] trains a neural network, which shares the same embedding layers with its generator, to detect watermarked texts. Similarly, one may develop a generic detector by training it to distinguish watermarked and non-watermarked texts. We explore this option in §5.
3.4 Watermark removal attacks
Attack Category Resource
Lowercasing Linguistic variation /
Contracting /
Expanding /
Misspelling Lexical editing Common misspellings
Typoing /
Synonymizing WordNet
Swapping /
Copy-pasting Text-mixing Non-watermarked text
Deep-paraphrasing Paraphrasing LLM-based paraphraser
Translating LLM-based translator
Black-box adversarial attack Generic detector
Table 3: A taxonomy of watermark removal attacks.
We also present a taxonomy of existing watermark removal attacks according to their underlying perturbation and required resources, as summarized in Table 3.
Linguistic variation attack. This class of attacks perturb the linguistic features of the watermarked text, without changing its semantics. We consider the set of perturbations in HELM[34], which simulate natural linguistic variations encountered in human interactions with text typing interfaces: i) Lowercasing converts all the words to lower-cases, which potentially affects the interpretation of proper nouns or emphases. ii) Contracting/expanding replaces phrases with their contracted/expanded forms (e.g., “I am” to “I’m” and vice versa), which may impact the tokenizer.
Lexical editing attack. This class of attacks modifies individual words, aiming to maintain the original text’s semantics. Specifically, we consider the following editing operations: i) Misspelling, similar to text-bugger[32], replaces words with their common misspellings (plural forms also considered); ii) Typoing replaces certain letters in a word with others; iii) Synonymizing replaces words with their synonyms using WordNet[43]; and iv) Swapping randomly exchanges the positions of two words within the text, which alters the text structure while potentially preserving the overall semantics.
Text-mixing attack. This class of attacks aims to “dilute” the watermark by mixing the watermarked text with non-watermarked text fragments. Specifically, the copy-pasting attack[27] embeds the watermarked text into the context of non-watermarked, human-written text. Note that the influence of non-watermarked text can be controlled by setting the fractions of watermarked and non-watermarked text fragments (i.e., the mixing weights).
Paraphrasing attack. This class of attacks relies on an additional LLM (i.e., paraphraser) to re-write the given watermarked text to evade the detector. For instance, a light paraphraser[52] (e.g., T5-based paraphraser[8]) can paraphrase the watermarked text sentence-by-sentence, while a more capable paraphraser (e.g., DIPPER[28]) paraphrases the watermarked text in one-shot, also enabling to control lexical diversity and token order diversity.
Similar to paraphrasing, the translating attack uses a translator LLM (e.g., Seamless [6]) to cycle the watermarked text through multiple languages (e.g., from English to French and back to English). This process can significantly alter the sentence structure and phrasing.
Black-box adversarial attack. In this class of attacks, we assume the adversary has access to a generic detector that is trained to distinguish watermarked and non-watermarked texts. The adversary then perturbs the watermarked text based on this surrogate detector. We explore this attack type in §6.
4 Platform
We implement representative watermarkers and watermark removal attacks on WaterPark, a platform for evaluating different watermarkers in a comprehensive and unified manner. With a modular design, WaterPark can be easily extended to incorporate new watermarkers and attacks.
4.1 Watermarkers and Attacks
Currently, WaterPark incorporates all the watermarkers in Table 2, including 8 one-bit and 2 multi-bit watermarkers. By default, the generator and detector of each watermarker are configured following its official documentation. In addition to the watermark-specific detector, we also implement a generic detector. Specifically, we approach the task as a binary classification problem. We gather watermarked and non-watermarked texts as positive and negative samples. The target LLM and the watermark generator jointly generate the positive samples, while the negative samples are non-watermarked texts generated by ChatGPT-3.5-Turbo (under the same prompt). We train a RoBERTa-based[39] binary classifier as the detector using the training data.
Meanwhile, we implement all the watermark removal attacks in Table 3. Specifically, the linguistic variation attacks follow the definitions of HELM[34]; the text-mixing attack is instantiated as the copy-pasting[27] attack; and the paraphrasing attack is instantiated as the Dipper[28] attack.
4.2 LLMs and Datasets
To factor out the influence of individual LLMs and datasets, WaterPark supports a range of LLMs and datasets.
Specifically, WaterPark integrates a variety of LLMs including OPT-1.3B[74], Llama2-7B[58], Mistral-7B[22], and Gemma-2B/7B[15]; it also includes a variety of datasets including: C4 is a cleaned version of Common Crawl’s web crawl corpus, which is used to train the T5 text-to-text Transformer models; HC3 (Human ChatGPT Comparison Corpus) consists of around 40K questions and their corresponding human/ChatGPT answers; LFQA (Long-Form QA)[13] is a long-form question-answering dataset that consists of around 24K question and answer pairs.
4.3 Metrics
WaterPark evaluates a given watermarker from multiple complementary aspects.
Effectiveness. At a high level, WaterPark evaluates the watermark detector’s accuracy in detecting watermarked texts mainly using two metrics: true positive rate (TPR) and false positive rate (FPR). TPR measures the fraction of watermarked texts detected as watermarked, while FPR measures the fraction of non-watermarked samples wrongly detected as watermarked. Formally, let $S_{+}$ and $S^{\prime}_{+}$ respectively be the sets of ground-truth and detected watermarked texts ($S_{-}$ and $S^{\prime}_{-}$ correspondingly).
$\text{TPR}=\frac{|S_{+}\cap S^{\prime}_{+}|}{|S_{+}|}\qquad\text{FPR}=\frac{|S_{-}\cap S^{\prime}_{+}|}{|S_{-}|}$ (9)
In WaterPark, we plot the receiver operating characteristic (ROC) curve that measures TPR against FPR across varying settings of detection thresholds. In particular, the area under the ROC curve (AUC) evaluates the overall effectiveness of each watermarker. Moreover, to compare two watermarkers under specific settings, we may also measure their TPRs under a fixed FPR (e.g., 1%).
Fidelity. To evaluate the impact of watermarking on text quality, WaterPark employs the following metrics to measure the difference between the original text $T$ and the watermarked text $\tilde{T}$.
WER (word error rate) measures the percentage of mismatched tokens between $\tilde{T}$ and $T$ relative to the total number of tokens in $T$. This is a lexical metric used to measure the fraction of tokens that have been modified.
BLEU[48] measures the lexical similarity of $T$ and $\tilde{T}$ by calculating the proportion of $n$-grams matched in $\tilde{T}$ and $T$.
BERTScore[75] measures the token-level similarity of $T$ and $\tilde{T}$ by leveraging the pre-trained contextual embeddings from BERT[10] and matching tokens in $T$ and $\tilde{T}$ using cosine similarity.
P-SP[65] evaluates the semantic similarity of $T$ and $\tilde{T}$ using the cosine similarity of their encodings. In particular, the encoding of $T$ (or $\tilde{T}$) is calculated by averaging the embeddings of its subword units generated by SentencePiece[30].
MAUVE[50] compares the distributions of $T$ and $\tilde{T}$ by computing an information divergence curve within a quantized embedding space. The area under this divergence curve provides a scalar summary of the trade-off between Type I ($T$ places high mass in areas where $\tilde{T}$ has low mass) and Type II (vice versa) errors.
Note that these metrics are also used to measure the impact of watermark removal attacks on the quality of the modified text $\tilde{T}^{\prime}$, relative to the watermarked text $\tilde{T}$.
Robustness. To evaluate a watermarker’s attack resilience, WaterPark measures the attack’s impact on the watermarking effectiveness. Specifically, let $\tilde{T}$ and $\tilde{T}^{\prime}$ respectively denote the watermarked text before and after the attack. WaterPark compares the detector’s TPR, FPR, and AUC with respect to $\tilde{T}$ and $\tilde{T}^{\prime}$. Intuitively, a smaller difference indicates that the watermarker is less attack-sensitive.
5 Evaluation
Leveraging WaterPark, we empirically assess representative LLM watermarkers, focusing on attack robustness and its relationships to other criteria. To our knowledge, this represents the most comprehensive evaluation of state-of-the-art watermarkers to date.
5.1 Experimental Setting
Our evaluation considers all the watermarkers in Table 2 and attacks in Table 3. We employ two LLMs OPT-1.3B and Llama2-7B-chat-hf to represent small and large models respectively. We use two datasets C4 and HC3 to simulate diverse tasks such as question-answering and text completion.
We implement and configure all watermarkers and attacks in accordance with their official documentation. Table 6 summarizes the default parameter settings. To validate our implementation, we assess their effectiveness and fidelity, as reported in §B and §C, corroborating results from existing literature. All experiments are conducted on a workstation running Nvidia RTX A6000 GPUs and 256GB of memory.
5.2 Robustness – Observational Study
Next, we systematically assess the resilience of existing LLM watermarkers against representative watermark removal attacks listed in Table 3. Our main findings are summarized in Table 4, demonstrating the impact of a watermarker’s various design choices on its robustness to various attacks.
Design Choice Attack Category
Linguistic var. Lexical edit. Text-mix. Paraphras.
Index-dep.
Text-dep.
Context-free
Dist.-shift
Dist.-reweight
Dist.-transform
Score-based
Diff.-based
Model-based
Table 4: Impact of a watermarker’s design choices on its attack robustness ( positive; negative; neutral)
Figure 2: Resilience of LLM watermarkers against linguistic variation attacks.
5.2.1 Linguistic Variation Attack
This class of attacks perturb the linguistic features of the watermarked text $\tilde{T}$, without changing its semantics. Here, we consider three attacks: lowercasing, contracting, and expanding, which are concretely defined in §3.4. We measure the TPRs of different watermarkers against such attacks (with their FPRs fixed as 1%). Further, to illustrate the impact of the watermarker design on its attack resilience. We group the results along three design choices: context dependency, sampling strategy, and detection strategy, as shown in Figure 2.
i) Overall, most watermarkers show strong resilience against linguistic variation attacks. For instance, TGRL reaches close to 100% TPRs under all three attacks. ii) UPV is the only watermarker marginally susceptible to such attacks. This can be explained by the fact that its neural network-based detector primarily depends on implicit textual features, which appear to be sensitive to changes in linguistic characteristics.
Remark 1 – Watermarkers that rely on explicit watermark signals tend to be robust against linguistic variation attacks.
5.2.2 Lexical Editing Attack
Figure 3: Resilience of LLM watermarkers against lexical editing attacks.
This class of attacks modifies individual words while maintaining the watermarked text’s semantics. Specifically, we consider four attacks in this category: synonymizing, misspelling, typoing, and swapping (see §3.4 for their concrete definitions). Similar to §5.2.1, we compare the TPRs of different watermarkers under these attacks (with their FPRs fixed as 1%). Figure 3 summarizes the results.
i) In terms of the context dependency design, text-dependent watermarkers (e.g., TRGL, UB, and GO) tend to be more vulnerable to lexical editing attacks, compared with index-dependent (e.g., RDF) and context-free (e.g., UG) watermarkers. Intuitively, as text-dependent watermarkers use previous tokens as the context for the next token in both the watermark generator and detector, the lexical editing thus causes a mismatch between the generator and detector. ii) Among the text-dependent watermarkers, it is observed that UB exhibits much higher vulnerability to such attacks, compared with the others. For example, its TPR drops near zero under the typoing and swapping attacks. This may be explained as follows. Recall that, to achieve unbiasedness, UB applies “hard” perturbation on the next-token distribution (e.g., by rejecting half of the vocabulary); thus the disruption to the previous tokens tends to cause a more significant mismatch between the generator and detector, compared with other watermarkers that employ “soft” perturbation (e.g., distribution shift and transform). iii) The typoing attack is particularly effective, as it may cause significant tokenization errors.
Remark 2 – Text-dependent, (hard) distribution-reweight watermarkers are more vulnerable to lexical editing attacks.
5.2.3 Text-Mixing Attack
Figure 4: Resilience of LLM watermarkers against text-mixing attacks.
This class of attacks “dilutes” the watermark by mixing the watermarked text with non-watermarked text fragments. Here, to evaluate the resilience of different watermarkers against text-mixing attacks, we use the copy-pasting attack[27] as the concrete attack, which embeds the watermarked text into the context of non-watermarked, human-written text (generated under the same prompt). We use CP-$n$-$m$ to denote the attack in which the modified text $T^{\prime}$ consists of $n$ segments of watermarked texts, each of length $m$of $|T^{\prime}|$, and the rest as non-watermarked text. We consider 4 different settings to assess the impact of the lengths and positions of the watermarked text segments. The results are shown in Figure 4.
i) Overall, most watermarkers experience significant TPR drops, especially under CP-$1$-$10$ that only preserves 10% of the watermarked text. ii) Among all the watermarkers, GO and RDF show significantly higher attack resilience. This can be attributed to their sampling strategy: both employ distribution transform, which generates the next token deterministically conditional on a given random permutation. Thus, GO and RDF tend to have stronger per-token signals than the other watermarkers that sample the next token from a given pool (e.g., green list). This observation is consistent with that in §B.2. iii) Meanwhile, UB and UPV are the most vulnerable to the copy-pasting attack, with close to zero TPRs under CP-$1$-$10$ and CP-$1$-$25$. This can be explained as follows. The model-assisted detector of UPV determines the given text as watermarked based on its aggregated features (rather than per-token statistics), while the injected non-watermarked segments may greatly disrupt such features. Meanwhile, UB applies hard perturbation on the next-token distribution (e.g., by rejecting half of the vocabulary); thus the disruption to the previous tokens causes a significant mismatch between the generator and detector.
Remark 3 – Distribution transform-based watermarkers show greater resilience to text-mixing attacks, while model-assisted, (hard) distribution-reweight watermarkers are more vulnerable.
Figure 5: Resilience of LLM watermarkers against paraphrasing attacks.
5.2.4 Paraphrasing Attack
This class of attacks employs an additional LLM (i.e., paraphraser) to re-write the given watermarked text $\tilde{T}$ (while preserving its semantics) to evade the detector. Here, we consider Dipper[28] as the paraphraser that rewrites $\tilde{T}$ in one shot. Specifically, we consider two settings of Dipper that respectively rewrite 20% and 40% of $\tilde{T}$. Further, we also consider the translating attack uses a translator model Seamless-m4t-v2-large[6] that first translates $\tilde{T}$ to French and then translates it back. Figure 5 compares the attack resilience of different watermarkers.
i) Both UPV and UB show higher vulnerability to the Dipper attack, compared with other watermarkers. For instance, the TPR of UPV drops to around zero under DP-$40$. This can be attributed to the same logic discussed in §5.2.3: the paraphrased text segments may greatly disrupt the aggregated textual features for UPV’s model-assisted detector, while the disruption to the previous tokens may cause a substantial mismatch between UB’s generator and detector, due to its rigid perturbation to the next-token distribution (e.g., rejecting half of the vocabulary). ii) Among the remaining watermarkers, RDF and UG are especially robust against the Dipper attack. This can be attributed to their index-dependent and context-free nature, which is less sensitive to the change of previous tokens than text-dependent watermarkers (e.g., TGRL, SIR, and GO). iii) Interestingly, RDF is more vulnerable to the translating attack than the Dipper attack. This can be explained by that RDF uses a detection strategy highly sensitive to the text length, while the translating attack can significantly shorten the text length.
Remark 4 – Index-dependent and context-free watermarkers tend to be more robust against paraphrasing attacks.
5.2.5 Fidelity Preservation
Recall that besides their attack effectiveness, another key metric for watermark removal attacks is whether they can preserve the quality of original texts. We thus compare the semantics of watermarked text $\tilde{T}$ and modified text $T^{\prime}$ using the metrics in §4. Figure 6 illustrates the quality preservation of different attacks on GO, with similar results on other watermarkers, with more results deferred to §D.3.
Figure 6: Quality preservation of different attacks on GO.
Observe that most attacks preserve the semantics of the watermarked text $\tilde{T}$ in the modified text $\tilde{T}^{\prime}$, as measured by BERTScore and P-SP scores. In comparison, the copy-pasting (CP) attack causes more significant text-quality degradation than other attacks, in that it may disrupt the orders of watermarked and non-watermarked segments and insert duplicate segments. Also, note that most attacks emphasize the semantic similarity between $\tilde{T}$ and $\tilde{T}^{\prime}$ rather than their lexical similarity (as measured by WER and BLEU scores).
5.3 Robustness – Causal Analysis
In addition to the observational studies, we further consider conducting causal analysis to understand the impact of individual design choices (e.g., samplers). However, this is challenging in our context because the various components of a watermarker are often highly interconnected and difficult to decouple. For example, the distribution-reweight strategy and the difference-based detection of UB are closely linked and cannot be easily replaced by other designs. To address this challenge, we select two watermarkers with their only difference in the design of one specific component (e.g., context dependency). The results are shown in Figure 7.
Figure 7: Resilience of LLM watermarkers against multi attacks. a) Different in Context Dependency: TGRL (text-dependent) and UG (context-free) b) Different in Generation Strategy: TGRL (distribution-shift) and GO (distribution-transform) c) Different in Detection Method: UPV (Model-based) and UPV-key (Score-based)
5.3.1 Context Dependency
We select TGRL and UG to represent text-dependent and context-free designs respectively. TGRL uses the previous $k$ tokens as a randomness seed to divide the vocabulary into red/green lists for the current token, whereas UG uses fixed red/green lists in generating all the tokens. In general, UG performs similarly to TGRL. However, when subjected to the Dipper attack at varying intensity levels, UG consistently outperforms TGRL. For instance, as we increase the attack intensity (e.g. DP-60-20, DP-60-40), which involves rearranging the text, UG maintains a relatively stable TPR with only minor reductions. In contrast, TGRL’s TPR significantly drops, particularly at DP-60-20. Recall that the Dipper attack extensively paraphrases and reorders the text, leading to substantial changes in consecutive tokens. Due to its text-dependent design, TGRL struggles to maintain consistency in random seeds between detection and generation. In contrast, UG’s context-free design avoids this issue. This aligns with the results in UG[76], showing that when subjected to paraphrasing attacks, UG consistently outperforms TGRL.
5.3.2 Generation Strategy
TGRL and GO are both context-dependent watermarkers. As it relies on distribution shift generation, TGRL tends to be less robust against copy-paste attacks. In contrast, although copy-paste attacks can significantly disrupt detection and dilute the watermark by inserting large amounts of text, due to its distribution transform design, GO can better maintain the watermark’s concentration. As a result, GO achieves a higher detection rate of watermarked text after such attacks.
5.3.3 Detection Method
UPV employs model-based detection to detect watermarks. Alternatively, it can also use score-based detection similar to TGRL. The original paper claims that model-based detection is more resistant to paraphrasing attacks than score-based detection. However, our results indicate that this advantage is not significant. Furthermore, score-based UPV demonstrates stronger robustness against weaker attacks (e.g., misspelling and swapping attacks). Using model-based detection introduces higher uncertainty compared to score-based detection and does not provide watermarkers with greater flexibility in resisting paraphrasing attacks.
Both RDF and GO utilize distribution-transform generation and rely on score-based detection, with RDF specifically employing edit score. RDF shows marginally higher robustness against lexical editing attacks with fixed token length compared to GO. However, it exhibits lower resilience against high-intensity text-mixing attacks (e.g., CP-3-10), where token length varies considerably.
6 Discussion
Next, we examine the current practices of operating watermarkers in adversarial environments and explore potential improvements.
6.1 Specific vs. Generic Detector
Figure 8: Detection of watermarked texts by watermarker-specific and generic detectors (‘1’ or ‘0’ indicate that the detector detects the given watermarked text as watermarked or non-watermarked).
For each watermarker, we mainly use its specific detector to detect watermarked texts. Here, we explore a generic, neural network-based detector as an alternative. To this end, we employ a pre-trained RoBERTa model and fine-tune it as a binary classifier using watermarked and non-watermarked texts. We use OPT-1.3B as the underlying LLM and C4 as the reference dataset. The watermarked text is generated by the LLM and the watermarker jointly, whereas the non-watermarked text is produced by ChatGPT-3.5-Turbo using the same prompt to mimic human response. Table 5 shows the TPRs of watermarkers with generic detectors (with FPRs fixed as 1%).
Watermarker TGRL RDF UB UG UPV SIR GO
TPR 0.9851 0.9812 0.9888 0.9997 0.9971 0.9962 0.9703
Table 5: TPRs of watermarkers with generic detectors (with FPRs fixed as 1%).
Next, we compare the attack resilience of watermarker-specific and generic detectors. For each watermarker, we apply the Dipper-$40$ attack and examine whether two detectors can effectively detect watermarked texts after the paraphrasing attack. Figure 8 depicts the confusion matrices of both detectors.
We have the following findings. i) The specific and generic detectors jointly achieve a high detection rate. Across all the watermarkers, the chance that both detectors fail to detect the watermarked texts (0-0) is below 0.07. ii) For RDF, the generic detector seems less effective than the specific detector, while for UB and UPV, the generic detector outperforms the specific detector by a large margin. Recall that UB and UPV are highly sensitive to the Dipper attack (see §5.2.4. Thus, employing a generic detector alongside a watermarker-specific detector can be an effective strategy for enhancing the security of vulnerable watermarkers. However, note that given the availability of generic detectors, it is also feasible for the adversary to leverage such detectors as an attack checker to adapt their attacks, which we will discuss in §6.2.3.
Remark 5 – Employing a generic detector alongside a watermark-specific detector improves the security of vulnerable watermarkers.
6.2 Advanced Attack
In addition to the previously mentioned simple attacks, we investigate the effectiveness of more advanced attacks.
6.2.1 Varying Attack Intensity
Figure 9: Resilience of watermarkers against increasingly intensive Dipper attacks.
One straightforward way to improve an attack’s effectiveness is to increase its intensity, potentially at the cost of other metrics (e.g., text quality). Here, we consider the Dipper attack[28] under varying intensity settings, denoted as DP-$l$-$o$, where lexical change ($l$) indicates that $l$of the given text is paraphrased and order change ($o$) indicates that $o$of the text is re-ordered. We compare the watermarkers’ resilience under varying attack intensity, as shown in Figure 9.
i) As expected, most watermarkers observe TPR drops as the attack intensity increases. ii) Among all the watermarkers, UG demonstrates the most consistent resilience under varying attack intensity. This can be attributed to its context-free design: the same perturbation is applied to the next-token distribution across all the tokens, which is thus immune to the change of previous tokens. iii) Compared with text-dependent watermarkers (e.g., TGRL and GO), an index-dependent watermarker (e.g., RDF) shows stronger resilience, especially under high attack intensity (e.g., DP-$60$-$20$), due to its weaker dependency on previous tokens. iv) UPV’s performance is somewhat inconsistent; it struggles with low-intensity attacks (e.g., DP-20) but shows resilience to high-strength ones (e.g., DP-60). This inconsistency can be attributed to UPV’s model-assisted detector and the inherent instability of its neural network, as confirmed by repeated experiments.
Remark 6 – A watermarker’s resilience to increasingly intensive attacks varies with its context dependency, roughly, context-free $>$ index-dependent $>$ text-dependent.
Figure 10: Resilience of watermarkers against individual (left) and combined (right) attacks.
6.2.2 Combining Simple Attacks
We first explore whether combining two attacks improves the attack’s effectiveness. Here, we only consider combining two simple attacks from the “weak” linguistic variation and lexical editing attacks. Figure 10 illustrates the effectiveness of such combined attacks.
We have a set of interesting observations. i) UPV and SIR, which demonstrate resilience against all simple attacks, are highly vulnerable to all the combined attacks. For instance, the TPR of SIR drastically drops to below 0.3 under the contracting+typoing attack. ii) UB, which is vulnerable to the typoing and swapping attacks, is consequently vulnerable to all the combined attacks that involve typoing or swapping. iii) TGRL and GO, which are robust against all the simple attacks (including typoing and swapping), show significant vulnerability to the typoing+swapping attack. This can be explained by that as typoing and swapping respectively disrupt the tokenization and token-indexing, their combination may substantially amplify such effects. iv) RDF and UG are especially robust against the combined attacks. This can be attributed to their “weaker” context dependencies, which is consistent with the findings in §5.2.4.
Remark 7 – Evaluating watermarkers’ robustness needs to consider not only individual attacks but also their combinations.
6.2.3 Adaptive Attack
Figure 11: Attacks leveraging surrogate detectors.
Given the availability of generic detectors, it is possible for the adversary to exploit such detectors to adapt their attacks. We consider an attack scenario as shown in Figure 11: the adversary performs a gradient-based adversarial attack[18] that iteratively modifies the watermarked text to evade the generic detector, and then forwards the modified text to the target, watermark-specific detector.
Figure 12: Detection of watermarked texts by watermarker-specific and generic detectors on gradient-based attacked samples (‘1’ or ‘0’ indicate that the detector detects the given watermarked text as watermarked or non-watermarked).
We evaluate this attack’s effectiveness using 500 watermarked texts from the C4 dataset. We use GBDA[18] as the adversarial attack and limit the number of perturbations to 100. The results are summarized in Figure 12. i) Leveraging the surrogate detector significantly improves the attack effectiveness: with the BERTScore between $\tilde{T}$ and $\tilde{T}^{\prime}$ is about 0.76, slightly lower than the Dipper attack, the detection rates of most watermarkers drop below 10%. ii) Although some samples do not evade the generic detector, they evade the specific detector successfully (e.g., TGRL, UG, and UB). iii) UG and RDF exhibit greater robustness than the other watermarkers. Specifically, for RDF, 42% of the samples evade the generic detector but are still detected by the RDF-specific detector. This superior robustness is likely due to their weaker context dependencies.
Remark 8 – Evaluating a watermarker’s robustness needs to consider attacks that leverage surrogate detectors.
Figure 13: Effectiveness of paraphrasing attacks with ChatGPT-3.5-Turbo as the paraphraser.
6.2.4 Leveraging Expert LLMs
We now explore the question of “what if the adversary has access to highly capable LLMs?” Specifically, we implement another paraphrasing attack that employs ChatGPT-3.5-Turbo as the paraphraser. For each watermarker, we randomly sample 100 watermarked texts that are successfully detected, and query the ChatGPT API with the following prompt:
[…]
and then forward the paraphrased text to the detector. To further evaluate the impact of the paraphrasing strength, we also measure the attack effectiveness under multiple rounds of paraphrasing as suggested in[72]. Notably, the attack described in [72] employs the T5 model to paraphrase the entire text up to 300 times. In contrast, our study limits paraphrasing to a maximum of 5 iterations. The difference between these threat models is due to the distinct objectives of the two studies: while [72] aims to understand the lower bound of watermark robustness, we focus on evaluating watermark robustness against practical, resource-limited adversaries. The results are summarized in Figure 13.
We have the following findings. i) The detection rates of all the watermarkers drop below 0.3 after one round of GPT paraphrasing, in which SIR and UB become nearly undetectable. This indicates that using highly capable LLMs to paraphrase watermarked texts is a dominant attack that effectively nullifies most existing watermarkers. However, as specified in our threat model (§2.3), access to highly capable LLMs may fall outside the scope of our robustness assessment, as their availability negates the need for watermark removal attacks. ii) For watermarkers that survive the first round of paraphrasing (e.g., UG and GO), their detection rates quickly drop below 0.15 as the adversary applies multiple rounds of paraphrasing. This observation corroborates the findings in[72].
Remark 9 – Paraphrasing using highly capable LLMs serves as a dominant attack to test watermarkers’ robustness lower-bounds.
6.3 Evaluation Guidelines
Next, we propose a set of guidelines for evaluating the robustness of LLM watermarkers. These guidelines incorporate our findings in §5 and §6, providing a minimal checklist to claim the robustness of an LLM watermarker.
LLMs and tasks. Our experiments show that the referenced watermarkers show varying robustness across different LLMs and datasets. We speculate that there exists an intricate interplay between the watermarking mechanism, the LLM’s capability, and the task’s complexity. We thus recommend experimenting on i) LLMs with varying capability (e.g., as measured by perplexity), ii) datasets for different tasks (e.g., summarization and question-answering), and iii) their combinations.
Attacks. Notably, using highly capable LLMs or applying computationally expensive rewriting can easily generate highly-quality, non-watermarked texts to evade detection; however, such attacks negate the need for watermark removal attacks in the first place. We thus recommend focusing on computationally efficient attacks such as linguistic variation, lexical editing, and lightweight paraphrasing, as well as their combinations, which reflects the risks of watermark removal attacks in practical settings.
Robustness. It is often critical to properly set the decision threshold for a watermarker (and also the attacks) to fully assess its robustness[41], which unfortunately is often missing in the original papers. To overcome this issue, we recommend i) measuring the overall effectiveness (TPR) in terms of ROC (across different threshold settings), ii) measuring the TPR under a fixed FPR (e.g., 0.01), and iii) considering varying attack intensity.
Fidelity. It is notoriously challenging to meaningfully measure the quality of text data[50]. We recommend employing a variety of metrics (e.g., BERTScore, P-SP, MAUVE) to comprehensively measure the quality retention of watermarkers as well as attacks. In addition, one may also leverage external tools (e.g., more advanced LLMs such as GPT-4) or human evaluation to provide a more accurate assessment if feasible.
7 Related Work
Next, we survey the relevant literature in the following categories: i) detection of LLM-generated texts, ii) LLM watermarking, iii) attacks on LLM watermarking, and iv) evaluation of LLM watermarkers.
Detection of LLM-generated texts. The advances in LLMs also give rise to their possible misuses[64, 63, 62, 61, 77, 16, 24, 9, 71, 33, 38, 23]. There is thus a pressing need for the capability of distinguishing LLM- and human-generated texts. Initial work attempts to either train classifiers using LLM- and human-generated texts[44] or to leverage intrinsic characteristics of LLM-generated texts (e.g., perplexity and variability in length, complexity, and information density)[42]. Yet, with LLMs becoming increasingly capable, the difference between LLM- and human-generated texts is narrowing, making such approaches less effective.
LLM watermarking In response, LLM watermarking emerges as a promising alternative, which instruments the LLM generative process with statistical signals that can be subsequently detected. The existing LLM watermarking techniques can be categorized based on the stages in which they are applied[37]: i) training-time watermarking[40, 57, 56, 55, 67], ii) watermarking during logit-generation[26, 76, 20, 35, 36, 12, 51, 14, 60, 69, 31, 17], iii) watermarking during token-sampling [29, 1, 19], and iv) post-generation watermarking[73, 70, 68, 45, 53, 2]. This study mainly focuses on training-free, pre-generation watermarking, which applies to any given LLMs and provides flexible control over multiple criteria (e.g., quality, effectiveness, and robustness).
The primary focus of the previous studies is to distinguish between human-written text and text generated by LLMs. In this paper, we assume that the attacker lacks access to non-watermarked text produced by recent LLM. Instead, the attacker can only utilize other LLMs to generate text that mimics human responses and get the watermarked texts from recent LLM. The generic detector, which is within the adversary’s capability, is specifically designed to distinguish between watermarked and non-watermarked texts(e.g., other LLMs’ texts or human writing texts).
Attacks on LLM watermarking. One critical property of an LLM watermarker is its robustness against potential attacks. A variety of attacks can be applied to LLM watermarking, ranging from removing the embedded watermark to uncovering the green/red lists. For instance, Dipper[28] is a widely used paraphrasing attack to evaluate the robustness of LLM watermarkers [26, 76] against watermark removal attacks; the watermark stealing attack[25] is proposed to identify green-list tokens and to replace them with red-list tokens, targeting TGRL[26] and UG[76] to further launch spoofing or removal attacks. This study primarily focuses on watermark removal attacks as they can target any LLM watermarkers and have profound implications in practice (e.g., disinformation, academic cheating, and automated phishing).
Evaluation of LLM watermarkers. As LLM watermarkers become prevalent, recent work attempts to benchmark the performance of various watermarkers. However, current studies either focus solely on the effectiveness of watermarking or have limited assessments of robustness. For instance, WaterBench[59] compares the effectiveness of TGRL[26] and UG[76] under varying hyper-parameter settings (e.g., prompt length); LLM-judger[54] uses GPT-3.5-Turbo as a judge to evaluate the effectiveness of RDF[29] and TGRL[26] and employs a binary classifier based on MLP to distinguish between watermarked and non-watermarked texts; MarkMyWords[49] compares the effectiveness of four watermarkers including TGRL[26], GO[1], RDF[29], and UW[5] and only evaluates the robustness of TGRL against watermark removal attacks.
To our best knowledge, this is the first study dedicated to the robustness of LLM watermarkers against watermark removal attacks. We aim to understand how different design choices impact watermarkers’ attack robustness and to identify best practices for operating watermarkers in adversarial environments.
8 Conclusion
In this paper, we systematize the existing LLM watermarkers and watermark removal attacks, mapping out the design space for various watermarking and attacking techniques. We then design and implement WaterPark, the first open-source platform devoted to assessing the attack robustness of LLM watermarkers in a unified and comprehensive manner. Leveraging WaterPark, we conduct a systematic evaluation of the robustness of existing watermarkers, addressing unresolved questions, revealing design trade-offs, and identifying opportunities for further improvement. Our findings shed light on the current LLM watermarking techniques, while WaterPark serves as a valuable benchmark aiding future research.
Appendix A Parameter Setting
Table 6 lists the default setting of the parameters of each watermarker in our evaluation. Note that $\gamma$ and $\delta$ are the gamma and delta used in the watermarker, and $n$ denotes the window size.
Watermarker Parameter Setting
TGRL $\gamma$ 0.25
$\delta$ 2.0
UG $\gamma$ 0.5
$\delta$ 2.0
CTWL $\delta$ 1.5
$n$ 10
message code length 20
encode ratio 10.0
message strategy vanilla
UPV $\gamma$ 0.5
$\delta$ 2.0
$n$ 3
bit number 16
layers 9
UB watermark type delta
SIR $\delta$ 1.0
$n$ 10
watermark type context
GO $n$ 3
RDF number of random sequences 50
Table 6: Default parameter setting of watermarkers.
Figure 14: Overall effectiveness of different watermarkers in generating and detecting watermarked texts: ROC of (a) OPT-C4, (b) OPT-HC3, (c) Llama2-C4, and (d) Llama2-HC3; TPR (with FPR fixed as 0.01) on (e) C4 and (f) HC3.
Appendix B Watermark Effectiveness
We first evaluate the effectiveness of different watermarkers. Following prior work[26, 49, 29], for each watermarker, we sample 1,000 prompts and use the LLM in combination with the watermarker to generate the watermarked texts; meanwhile, we select human responses to the same prompts as the non-watermarked texts. We then measure the accuracy of the watermarker’s detector in distinguishing the watermarked and non-watermarked texts.
B.1 Overall Effectiveness
We measure the overall effectiveness of each method through the lens of the ROC curve. Figure 14 (a-d) summarizes the overall effectiveness of existing watermarkers across different models and datasets. We have the following interesting observations.
Most watermarkers are highly effective in generating and subsequently detecting watermarked texts on OPT-1.3B as shown in Figure 14 (a-b). For instance, RDF, UB, and GO all attain AUC scores above 0.99 over both C4 and HC3. Recall that C4 and HC3 represent the text completion and question-answering tasks respectively. The observation indicates that most watermarkers tend to be highly effective for relatively less capable LLMs such as OPT-1.3B, while the concrete dataset/task have a limited impact on their performance. We further validate this hypothesis under the setting of a fixed FPR. As shown in Figure 14 (e-f), we fix the FPRs of all the methods to be 0.01 and measure their TPRs. Observe that all the methods achieve above 0.9 TPR, with a marginal difference across C4 and HC3.
Meanwhile, most methods observe marginal performance drops on Llama2-7B, as shown in Figure 14 (c-d). For instance, compared with its performance on OPT-1.3B, the AUC of SIR drops by 0.11 and 0.08 on C4 and HC3 respectively. This observation aligns with previous research [29], indicating that watermarkers are more effective on OPT compared with Llama2. This phenomenon can be partly understood through the following explanation. In contrast of less capable LLMs (e.g., OPT-1.3B), Llama2-7B typically produces texts of lower perplexity. Since most watermarkers inject watermarks by slightly altering the next-token distribution, the lower perplexity in Llama2’s outputs hampers the effectiveness of such perturbations. Moreover, it is observed that the performance of various methods varies significantly across different datasets. For instance, UG’s AUC differs by 0.28 between C4 and HC3, while UB’s AUC differs by 0.11. This observation is further supported by the TPR measures at fixed FPRs (fixed as 1%), as shown in Figure 14 (e-f). Our findings suggest that the concrete dataset/task tends to have a larger impact on watermarkers over more capable LLMs.
Observation 1 – Ensuring watermarkers achieve the same quality and effectiveness as reported in their works and existing watermarkers tend to be more effective on LLMs with higher perplexity.
B.2 Impact of Text Length
Figure 15: TPRs of watermarkers with respect to text length (with FPRs fixed as 1%).
We evaluate how the (non-)watermarked text length (i.e., the number of tokens) impacts the performance of different methods. Specifically, we measure the TPR of each method with its FPR fixed as 1%. In the following, we set OPT-1.3B and C4 as the default LLM and dataset. Figure 15 summarizes the results.
Observe that as expected, the TPRs of all the methods improve as the text length grows from 1 to 200 tokens. As the text length exceeds 100 tokens, most methods reach TPRs close to 100%. Meanwhile, different methods show varying sensitivity to the text length. For instance, GO and RDF attain 100% TPRs with only 20 tokens, while UG reaches only around 50% TPR under the same setting. This can be explained as follows. Both GO and RDF use distribution transform-based samplers, which, conditional on given randomness (e.g., random permutation), generate the next token deterministically. Meanwhile, other methods randomly sample the next token from a given pool (e.g., green lists). Thus, GO and RDF tend to have stronger signals per token for watermark detection.
Observation 2 – Different watermarkers have varying sensitivity to text length; distribution transform-based watermarkers are the least sensitive.
B.3 Impact of Temperature
The temperature $\tau$ is a key parameter that affects a watermarker’s generative dynamics: intuitively, a higher $\tau$ makes the sampling over the next-token distribution $\tilde{p}(x_{t}|x_{<t})$ more random. Here we evaluate how the setting of $\tau$ in each watermaker’s generator may impact its effectiveness. Note that, unlike other watermarkers, RDF and GO do not generate the next-token distribution $\tilde{p}(x_{t}|x_{<t})$ explicitly, we thus exclude them from the evaluation.
Figure 16: TPRs of watermarkers with respect to the temperature setting (with FPRs fixed as 1%).
Figure 16 compares how the TPRs of different watermarkers vary with the setting of $\tau=$ 0.7, 1.0, and 1.3 (with FPRs fixed as 1%). Observe that the performance of most watermarkers marginally improves with $\tau$, which corroborates prior work[49]. For instance, the TPR of TGRL increases by about 0.05 as $\tau$ varies from 0.7 to 1.3. Interestingly, in contrast, the TPR of UPV decreases as $\tau$ grows. This can be explained as follows. UPV employs a neural network as the detector that depends on general textual features, which tends to be more sensitive to increasing randomness, compared with other watermarkers that rely on specific watermark signals (e.g., green/red-listed tokens).
Observation 3 – The effectiveness of watermarkers that rely on explicit watermark signals improves with the temperature.
Appendix C Watermark Fidelity
Figure 17: Fidelity preservation of different watermarkers.
We evaluate the impact of different watermarkers on the text quality. We compare the original text $T$ and watermarked text $\tilde{T}$ using the metrics (detailed in §4) of BERTScore[75], P-SP[65], and MAUVE[50]. The results are summarized in Figure 17.
i) A majority of watermarkers well preserve the semantics of original texts, as indicated by their high BERTScore and MAUVE scores. Note that the P-SP scores of all the watermarkers are relatively lower than their BERTScore and MAUVE scores. This is due to their different emphases: P-SP measures the average similarity between the tokens in $T$ and $\tilde{T}$, while BERTScore calculates the maximum similarity between the tokens in $T$ and $\tilde{T}$. ii) Meanwhile, RDF, GO, and UB are less effective in preserving the quality of original texts, which can be attributed to their additional constraints of indistinguishability: the expectation of the watermarker’s next-token distribution is identical to the LLM’s next-token distribution (i.e., indistinguishability). This observation suggest that there exists an inherent trade-off between the desiderata of quality and indistinguishability.
Observation 4 – There exists a fundamental trade-off between the watermarker’s fidelity and indistinguishability.
Appendix D Additional Results
D.1 Multi-bit Watermarking
While our study focuses on one-bit watermarkers, for completeness, we also evaluate CTWL[60], a multi-bit watermarker. In contrast of one-bit watermarkers that encode only a single bit of information (i.e., whether a given text is watermarked), a multi-bit watermarker can encode multiple bits of information into the watermarked text, such as the generating model, the date of generation, and other details. However, despite its larger information capacity, we find that a multi-bit watermarker is typically less robust compared to one-bit watermarkers, as illustrated in Figure 18 and  19.
Figure 18: TPRs of CTWL (with FPRs fixed as 0.01) on different LLMs (OPT and Llama2) and datasets (C4 and HC3).
Figure 19: TPRs of CTWL (with FPRs fixed as 0.01) against various attacks.
The test results in the basic encoding and detection scenario highlight the sensitivity of CTWL to different language models. Although it performs well with the OPT model across both C4 and HC3 datasets, its TPR on Llama2 is extremely low, and it becomes completely ineffective when tested on Llama2 using the HC3 dataset. This is due to the lower model perplexity of Llama2. We conduct additional experiments to compare the model perplexity on the same WikiText dataset, and the results show that Llama2 has a perplexity about 6.15, while OPT has about 12.43. Lower model perplexity results to larger fluctuations in the logits produced by the model, makes it more challenging for watermark injection (e.g., increase smaller logits to exceed larger ones).
When facing the attacks, CTWL is more vulnerable than one-bit methods. It is particularly vulnerable to the copy-pasting attack, which can nearly disable the method as the TPR drops to near zero. Additionally, CTWL is highly susceptible to typoing, swapping, translating, lowercasing, and Dipper attacks, which generally do not affect many one-bit methods as severely. Despite the capacity of multi-bit methods to embed more information, their high sensitivity to language model variations and various attacks is a crucial limitation that needs to be addressed in future research.
D.2 Robustness of Watermarkers on Llama2
Our evaluation in §5 mainly uses OPT-1.3B as the underlying LLM. Here, we present the results on Llama2-7B, a more capable LLM.
D.2.1 Linguistic Variation Attack
Comparing with the OPT results, most of the watermarkers show TPR drops on Llama2. One exception is GO, which still maintains resilience across all three attacks.
Figure 20: Resilience of LLM watermarkers on Llama2 against linguistic variation attacks.
D.2.2 Lexical Editing Attack
We have the following observation in Llama2: i) Context-free watermarkers (e.g., UG) show significantly higher attack resilience against swapping attacks. This may be due to the fact that the attack alters the word and sentence order, which affects the detection of both text-dependent and index-dependent watermarkers. ii) Similar to OPT, text-dependent, (hard) distribution-reweighting watermarkers tend to be more vulnerable to lexical editing attacks. iii) The typoing and swapping attacks show significantly higher attack effectiveness on Llama2, similar to our observations on OPT.
Figure 21: Resilience of LLM watermarkers on Llama2 against lexical editing attacks.
D.2.3 Text-Mixing Attacks
With the exception of watermarkers based on distribution transform (e.g., RDF and GO), which shows greater resilience to text-mixing attacks, the other methods are virtually undetectable, which is consistent with the observations in OPT
Figure 22: Resilience of LLM watermarkers on Llama2 against text-mixing attacks.
D.2.4 Paraphrasing Attack
All methods show a significant TPR decrease, only GO presents a remarkable resilience. RDF performs well in OPT and does not present a significant resilience on Llama2.
Figure 23: Resilience of LLM watermarkers on Llama2 against paraphrasing attacks.
D.3 Fidelity Preservation
Figure 24 illustrates the quality preservation of different attacks on GO, with similar results on other watermarkers,
Figure 24: Quality preservation of different attacks on GO.